# Speculative Ensemble: Fast Large Language Model Ensemble via Speculation

### v1

@@ B.2 Speculative Ensemble Framework
At this stage, $x_{1},x_{2},x_{3}$ have been scored by all models, allowing the ensemble distributions $r_{1}(x),r_{2}(x),r_{3}(x)$ to be computed and used for verification, as illustrated in step 3. Assuming that $x_{1},x_{2},x_{3}$ are all accepted, their corresponding probability distributions are no longer needed and are discarded, as shown in step 4. Subsequently, in step 5, $\mathcal{M}_{1}$ is invoked again to score $x_{4},x_{5},x_{6}$ and generate new proposal tokens: $x_{7},x_{8},x_{9}$.
@@
Figure 7: The sketch of speculative ensemble framework in three-model ensemble scenario. The colored boxes represent the stored probability distributions, while the grey boxes represent the cleared ones. Each invocation involves scoring the current proposal tokens and generating a bonus token. The proposal length for model $\mathcal{M}_{1},\mathcal{M}_{2},\mathcal{M}_{3}$ is 3, 2, 1, respectively.
Algorithm 2 The speculative ensemble framework. Propose employs model $\mathcal{M}$ to take the current token sequence $S$ and a constant $\gamma$ as inputs and generates $\gamma$ tokens $T$. Score employs model $\mathcal{M}$ to score a sequence to obtain sequence probabilities $P$, a bonus token $t$ and its probability $p$ Verify examines $T$ to decide whether it should be accepted according to ensemble probability distribution.
@@
           $p^{\prime}\leftarrow\mathcal{E}(p_{11},\dots,p_{n1})$$\triangleright$ $p_{ij}$ represents the $j$-th probability of $C_{i}$
@@ Appendix C Additional Results
C.1 Speculative Ensemble for Quality-Speed Tradeoff
As outlined in 3.2, creating a weighted ensemble of the proposal and target models in Speculative Ensemble (SE) offers a way to balance the tradeoff between quality and speed. We conducted experiments with the Llama-3 model pair on four datasets, adjusting $\lambda$ from 0.1 to 0.9. As shown in Figure 8, increasing $\lambda$ leads to a steady improvement in inference speed but a gradual decline in performance, allowing users to choose a tradeoff that best suits their needs.
@@ C.1 Speculative Ensemble for Quality-Speed Tradeoff
Figure 8: Speculative Ensemble for quality-speed tradeoff
@@ C.2 Speedup on OPT Model Pair
CD-SE 1.69x 1.76x 2.16x 1.85x

### v2

@@ B.2 CoS Framework
At this stage, $x_{1},x_{2},x_{3}$ have been scored by all models, allowing the combined distributions $r_{1}(x),r_{2}(x),r_{3}(x)$ to be computed and used for verification, as illustrated in step 3. Assuming that $x_{1},x_{2},x_{3}$ are all accepted, their corresponding probability distributions are no longer needed and are discarded, as shown in step 4. Subsequently, in step 5, $\mathcal{M}_{1}$ is invoked again to score $x_{4},x_{5},x_{6}$ and generate new proposal tokens: $x_{7},x_{8},x_{9}$.
@@
Figure 7: The sketch of CoS framework in three-model scenario. The colored boxes represent the stored probability distributions, while the grey boxes represent the cleared ones. Each invocation involves scoring the current proposal tokens and generating a bonus token. The proposal length for model $\mathcal{M}_{1},\mathcal{M}_{2},\mathcal{M}_{3}$ is 3, 2, 1, respectively.
Algorithm 2 The CoS framework. Propose employs model $\mathcal{M}$ to take the current token sequence $S$ and a constant $\gamma$ as inputs and generates $\gamma$ tokens $T$. Score employs model $\mathcal{M}$ to score a sequence to obtain sequence probabilities $P$, a bonus token $t$ and its probability $p$ Verify examines $T$ to decide whether it should be accepted according to combined probability distribution.
@@
           $p^{\prime}\leftarrow\mathcal{C}(p_{11},\dots,p_{n1})$$\triangleright$ $p_{ij}$ represents the $j$-th probability of $C_{i}$
@@ Appendix C Additional Results
C.1 CoS for Quality-Speed Tradeoff
As outlined in 3.2, creating a weighted ensemble of the proposal and target models in CoS offers a way to balance the tradeoff between quality and speed. We conducted experiments with the Llama-3 model pair on four datasets, adjusting $\lambda$ from 0.1 to 0.9. As shown in Figure 8, increasing $\lambda$ leads to a steady improvement in inference speed but a gradual decline in performance, allowing users to choose a tradeoff that best suits their needs.
@@ C.1 CoS for Quality-Speed Tradeoff
Figure 8: CoS for quality-speed tradeoff
@@ C.2 Speedup on OPT Model Pair
CD-CoS 1.69x 1.76x 2.16x 1.85x
C.3 The Raw Speed
The results are shown in Table 5 and Table 6.
Table 5: The raw speed of each method under WE setting. The table reports the average number of tokens generated per second. Models are of comparable sizes.
Model Pair Method HumanEval GSM8K MMLU CNNDM
Llama-Vicuna WE 22.617 22.054 20.459 20.782
WE-SD 28.723 26.685 24.346 23.899
WE-CoS 35.734 33.522 28.847 30.341
Qwen-3b WE 49.624 49.025 47.764 46.966
WE-SD 56.075 51.966 52.062 50.723
WE-CoS 80.390 74.518 67.824 64.813
Qwen-1.5b WE 82.584 77.941 79.631 76.625
WE-SD 91.668 88.073 86.001 84.287
WE-CoS 128.831 113.793 106.705 103.443
Qwen-1.5b WE 57.048 53.869 53.742 56.286
WE-SD 55.050 51.082 52.666 51.927
WE-CoS 105.485 68.287 74.426 85.980
Table 6: The raw speed of each method under CD setting. Models are of different sizes.
Model Pair T Method HumanEval GSM8K MMLU CNNDM
Llama-3 0 CD 39.387 38.735 38.969 38.100
CD-SD 80.349 70.110 59.233 60.198
CD-CoS 87.833 77.470 68.975 61.341
1 CD 38.981 38.585 39.028 38.124
CD-SD 60.421 46.688 46.834 40.793
CD-CoS 64.319 55.562 51.127 44.986
Llama-2 0 CD 49.314 48.129 49.700 46.215
CD-SD 56.711 77.969 53.676 42.980
CD-CoS 62.136 79.413 83.496 60.080
1 CD 47.325 47.015 47.492 43.512
CD-SD 44.486 54.537 58.415 47.863
CD-CoS 54.424 56.418 65.064 48.298
OPT 0 CD 23.525 23.525 19.576 19.692
CD-SD 23.211 24.701 28.776 27.568
CD-CoS 78.487 61.400 66.949 77.783
1 CD 23.934 23.422 19.492 19.671
CD-SD 35.182 36.304 41.128 36.391
CD-CoS 40.448 41.222 42.102 36.391
C.4 Ablation study on Alternate Proposal Framework
The results are shown in Table 7 and Table 8. Note that CD-CoS without APF is equivalent to SCD.
Table 7: Ablation study on alternate proposal framework (APF) under weighted ensemble (WE) setting ($\lambda=0.5$). Models are of comparable sizes.
Model Pair Method HumanEval GSM8K
Llama-Vicuna w/o APF 1.46x 1.39x
WE-CoS 1.58x 1.52x
Qwen-3b w/o APF 1.40x 1.32x
WE-CoS 1.62x 1.52x
Table 8: Ablation study on alternate proposal framework (APF) under contrastive decoding (CD) setting ($\mu=0.1$). Models are of different sizes.
Model Pair T Method HumanEval GSM8K
Llama-3 0 w/o APF 1.98x 1.72x
CD-CoS 2.23x 2.00x
1 w/o APF 1.41x 1.24x
CD-CoS 1.65x 1.44x
Llama-2 0 w/o APF 1.18x 1.59x
CD-CoS 1.26x 1.65x
1 w/o APF 0.89x 1.21x
CD-CoS 1.15x 1.20x
C.5 Speedup on NPUs
The results are shown in Table 9 and Table 10.
Table 9: The raw speed of each method under both WE and CD settings using the Ascend 910B3 NPU.
Model Pair Method HumanEval GSM8K MMLU CNNDM
Llama-Vicuna WE 7.60 5.83 2.60 2.35
WE-SD 13.59 8.73 3.81 3.55
WE-CoS 14.14 9.98 4.32 4.06
Llama-2 CD 11.98 10.63 4.96 4.51
CD-SD 13.00 16.28 7.098 4.83
CD-CoS 16.71 16.30 7.54 5.23
Table 10: The speedup ratio of each method under both WE and CD settings using the Ascend 910B3 NPU.
Model Pair Method HumanEval GSM8K MMLU CNNDM
Llama-Vicuna WE 1.00x 1.00x 1.00x 1.00x
WE-SD 1.79x 1.50x 1.47x 1.51x
WE-CoS 1.86x 1.71x 1.66x 1.73x
Llama-2 CD 1.00x 1.00x 1.00x 1.00x
CD-SD 1.09x 1.53x 1.43x 1.07x
CD-CoS 1.39x 1.53x 1.52x 1.16x